OpenStudy (anonymous):

Multiply (3 + 2i)(-1 + 4i).

OpenStudy (anonymous):

is this a complex equation the 2i and 4i part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok give me a minute it been a while doing this by hand.....easier with a calculator

OpenStudy (anonymous):

ok so this is what you have \[(3+2i)(-1+4i)\rightarrow-3+12i-2i-8\rightarrow-11+10i\] when you use the foil method and you multiply the 2i and 4i together you get \[8i^2=-8\] remember that. the negative sign will always apply when you multiply two complex roots together so your final answer will be \[-11+10i\] just like i got above

OpenStudy (anonymous):

Solve using the Quadratic Formula. 2x2 + 3x - 6 = 0

OpenStudy (anonymous):

do you know the quadratic eq?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok so you have \[ax^2+bx+c=0\] and from the problem you are solving from you can tell what your a,b, and c value is now the quadratic equation is \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] use that formula to solve for your x values. you will have 2 x values

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome anytime